Us the method of substitution by solving:
x-y=13
5x-5y=15
Can you show me the work so maybe I can get it?
thank you, thank you!

Answer by sachi(548)

x-y=13.........1
5x-5y=15.......2
from 1
x=y+13
substituting this in eqn 2
5(y+13)-5y=15
or 5y+65-5y=15
or 65=15
so the question is wrong